The radius of the circle having centre at (2,1) and whoseone of the chord is diameter of the circle x2 + y2 – 2x – 6y + 6 =0 is??Sir please explain!!

Grade:11

1 Answers

Sunil Raikwar
9 years ago
let the radius of circle be r then equation of circle
(x-2)2+(y-1)2=r2
x2+y2-4x-2y+5-r2=0...................1
x2+y2-2x-6y+6=0........................2
common chord of circle 1 & 2
s1-s2=0
2x-4y+1+r2=0
this passes through (1,3)
2-12+1+r2=0
r2=9
r=3